You are given the following information about an
economy:
Gross private domestic investment = 40
Government purchases of goods and services = 30
Gross national product (GNP) = 200
Current account balance = -20
Taxes = 60
Government transfer payments to the domestic private
sector = 25
Interest payments from the government to the domestic
private sector = 15 (Assume all interest payments by
the government go to domestic households.)
Factor income received from rest of world = 7
Factor payments made to rest of world = 9
Find the following, assuming that government investment
is zero:
a. Consumption
b. Net exports
c. GDP
d. Net factor payments from abroad
e. Private saving
f. Government saving
g. National saving
1
Expert's answer
2018-08-10T06:56:08-0400
a. Consumption: C = Y(GDP) – I – G -NX
С = 202 – 40 -30 – (-18) = 150
b. Net exports: NX = CA – NFP = -20 - (7-9) = -18
c. GDP = GNP – NFP
GDP = 200 - (7-9) = 202
d. Net factor payments from abroad: NFP = 7 – 9 = 2
e. Private saving: Spvt = Y – C = Y + NFP – T +TR + INT – C
Spvt = 202 -2 – 60 +25 + 15 -150 = 30
f. Government saving: Sgovt= T – TR – INT – G
Sgovt = 60 – 25 – (-15) – 30 = -10
g. National saving: S = Spvt + Sgovt
S = 30 + -10 = 20